# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ 2.3 Online Minibatch SGD
$(2C_{1}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}\preceq\Sigma(\mathbf{v})\preceq(2C_{2}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\mathbf{W}\mathbf{H}\mathbf{W}^{\top}.$
Noting $\nabla^{2}\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$ and $\mathcal{R}(\mathbf{v})=\mathcal{E}(\mathbf{v})+\frac{1}{2}\sigma^{2}$, this lemma means $\Sigma(\mathbf{v})\eqsim\mathcal{R}(\mathbf{v})\nabla^{2}\mathcal{R}(\mathbf{v}).$ That is, the gradient noise scales proportionally with the population risk and aligns with the local curvature. Notably, the noise has two distinct sources: (i) the fit-dependent term $\mathcal{E}(\mathbf{v})$, which arises purely from minibatching and persists even in the absence of label noise; (ii) the $\sigma^{2}$ term, which captures the contribution from label noise. This anisotropic structure of SGD noise – scaling with risk and shaped by curvature – has also been observed in prior work (wu2022alignment, ; wu2023implicit, ).
3 The Intrinsic-Time Stochastic Differential Equations
We will leverage the view of SDE modeling of SGD, a framework that has proven effective in characterizing the dynamical behavior of SGD (li2017stochastic, ; li2019stochastic, ; li2020reconciling, ; li2021validity, ; pesme2021implicit, ; li2022what, ). By moving to continuous time (su2016differential, ), the dynamics become both more tractable and more interpretable.
Fix a discretization step size $h$ and let $\varphi_{k}=\eta_{k}/h$ for all $k\in\mathbb{N}$. Then, (4) can be written as $\mathbf{v}_{k}=\mathbf{v}_{k-1}-\varphi_{k}\nabla\mathcal{R}(\mathbf{v}_{k-1})h-\varphi_{k}h\bm{\xi}_{k}.$ For small $h$, this recursion is approximated by the Itô-type time-inhomogeneous SDE li2019stochastic ; orvieto2019continuous ; ankirchner2024comparison :
$\mathop{}\!\mathrm{d}\bar{\mathbf{v}}_{\tau}=-\varphi(\tau)\nabla\mathcal{R}(\bar{\mathbf{v}}_{\tau})\mathop{}\!\mathrm{d}\tau+\varphi(\tau)\sqrt{\frac{h}{b(\tau)}\Sigma(\bar{\mathbf{v}}_{\tau})}\mathop{}\!\mathrm{d}\mathbf{B}_{\tau},$ (5)
where $\mathbf{B}_{\tau}\in\mathbb{R}^{M}$ denotes the $M$-dimensional Brownian motion, and
  • $\varphi\in C(\mathbb{R}_{\geqslant 0})$ is the continuous LRS function with $\varphi(kh)=\eta_{k}/h$ for all $k\in\mathbb{N}$;
  • $b\in C(\mathbb{R}_{\geqslant 0})$ is the continuous batch-size schedule function with $b(kh)=B_{k}$ for all $k\in\mathbb{N}$.
One can explain $\tau$ as the continuous steps or physical time.
The intrinsic-time SDE.
To isolate the effect of LRS, we introduce the concept of intrinsic time
$t=T(\tau):=\int_{0}^{\tau}\varphi(r)\mathop{}\!\mathrm{d}r,$
which counts the intrinsic time accumulated by training for $\tau$ steps.
Let $\bm{\nu}_{t}=\bar{\mathbf{v}}_{T^{-1}(t)}$ and applying the Øksendal’s time change formula (oksendal2003stochastic, ) to (5) gives the intrinsic-time SDE:
$\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\nabla\mathcal{R}(\bm{\nu}_{t})\mathop{}\!\mathrm{d}t+\sqrt{\gamma_{\varphi,b}(t)\Sigma(\bm{\nu}_{t})}\mathop{}\!\mathrm{d}\mathbf{B}_{t}\,\,\text{ with }\,\,\gamma_{\varphi,b}(t)=\frac{h\varphi(T^{-1}(t))}{b(T^{-1}(t))}.$ (6)
Note that $\gamma_{\varphi,b}(t)$ captures the effective gradient noise intensity, modulated by both the learning rate and batch size scheduling. Small learning rates and large batch sizes make this term smaller. Alternatively, the discrete SGD update (4) can be intuitively viewed as an “Euler–Maruyama discretization” of the SDE (6) on the non-uniform grid $\{t_{k}=\sum_{j=0}^{k}\eta_{j}\}_{k\in\mathbb{N}}$, as given by
$\mathbf{v}_{k}-\mathbf{v}_{k-1}=-\nabla\mathcal{R}(\mathbf{v}_{k-1})(t_{k}-t_{k-1})-\sqrt{\eta_{k}}\sqrt{t_{k}-t_{k-1}}\bm{\xi}_{k}.$
4 Functional Scaling Laws
Before stating our results, we introduce two key functions:
$e(t):=\int_{M^{-\beta}}^{1}u^{s-1}e^{-2ut}\mathop{}\!\mathrm{d}u,\quad\mathcal{K}(t):=\int_{M^{-\beta}}^{1}u^{1-\frac{1}{\beta}}e^{-2ut}\mathop{}\!\mathrm{d}u.$
The function $e(\cdot)$ characterizes the risk convergence of full-batch GD in terms of intrinsic time. The function $\mathcal{K}(\cdot)$, termed the forgetting kernel, is a decreasing function that captures how the model gradually forgets the noise memorized in earlier training. When $M=\infty$, $e(t)\eqsim t^{-s},\mathcal{K}(t)\eqsim t^{-(2-1/\beta)}$. For finite $M$, see Figure 1(left) for an illustration.
Figure 1: (left) The functions $e(\cdot)$ and $\mathcal{K}(\cdot)$. (middle, right) FSL accurately captures the average-risk dynamics of discrete SGD. Shaded regions indicate the mean over 200 independent runs, while solid lines represent FSL predictions. Results are shown for three LRSs: cosine, WSD-like, and a non-standard cyclic schedule. Additional experimental details and results for various values of $s$ and $\beta$ are provided in Appendix B.2.
Our main result is stated in the following theorem, whose proof can be found in Appendix C.
Theorem 4.1 (Intrinsic-time Functional Scaling Law (FSL)).
Suppose Assumptions 2.2 and 2.3 hold. Then, for the intrinsic-time SDE with top-$M$ features and $s\in(0,2-1/\beta]$, the following FSL holds for sufficiently large $M$ and $t$:
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\underbrace{\frac{1}{2}\sigma^{2}}_{\mathrm{irreducible}}\eqsim\underbrace{\frac{1}{M^{s\beta}}}_{\mathrm{approximation}}+\underbrace{\frac{1}{t^{s}}}_{\mathrm{full-batch\,GD}}+\underbrace{\int_{0}^{t}\mathcal{K}(t-r)(e(r)+\sigma^{2})\gamma_{\varphi,b}(r)\mathop{}\!\mathrm{d}r}_{\mathrm{noise\,term}\,\,\mathcal{N}(\varphi,b)}$ (7)
For the random feature case, the same FSL holds for $s\in[0,1]$ with probability at least $1-e^{-\Omega(M)}$ with respect to the randomness of $\mathbf{W}$.
This theorem establishes a scaling law for the expected population risk of the intrinsic-time SDE (6), where the influence of scheduling learning rate and batch size is explicitly described via the convolution-type noise functional: $(\varphi,b)\mapsto\mathcal{N}(\varphi,b)$. Notably, each term of our FSL is interpretable:
  • The irreducible risk ($\frac{1}{2}\sigma^{2}$) is due to the label noise.
  • The model-size scaling (approximation error): The corresponding scaling exponent depends only on the intrinsic difficulty $s\beta$.
  • The intrinsic-time scaling (full-batch GD term): The scaling exponent depends only on the relative difficulty $s$. Thus, for a fixed target, choosing a higher-capacity model (smaller $\beta$) make it converge faster.
  • The noise term $\mathcal{N}(\varphi,b)$ exhibits a simple convolution form, capturing the risk arising from fitting the noise injected by SGD during training. Specifically, $(e(r)+\sigma^{2})\gamma_{\varphi,b}(r)$ denotes the magnitude of gradient noise, adjusted by the learning-rate and batch-size schedules, at time $r$. More precisely, $e(r)\gamma_{\varphi,b}(r)$ represents the noise arising from mini-batch approximation, while $\sigma^{2}\gamma_{\varphi,b}(r)$ corresponds to the noise introduced by label noise. The forgetting kernel $\mathcal{K}(t-r)$ represents how much fitting noise at time $r$ is forgotten by time $t$.
FSL captures the risk dynamics of SGD.
Figure 1 (middle, right) compares the average risk dynamics of discrete SGD with the predictions of FSL. Despite being derived in continuous time, FSL closely captures the empirical behavior of SGD across different LRSs, task difficulties, and model capacities.
The trade-off between increasing intrinsic time and reducing noise injection.
The adjustment factor $\gamma_{\varphi,b}(\cdot)$ encodes the influence of learning-rate and batch-size schedules for noise injection. Smaller learning rates or larger batches reduce this factor, but they also shrink the intrinsic training time. This time budget is crucial in two ways: (i) the full‑batch GD term needs it to drive the risk down, and (ii) the forgetting kernel needs it to wash out noise memorized in the early iterations. A good schedule choice should therefore strike a compromise – suppressing injected noise while still preserving enough intrinsic time to accomplish both objectives.
The emergence of power laws.
We illustrate how the power law emerges in our setting from a multi-task learning viewpoint. For brevity, consider the case of the top-$M$ features and an infinitesimal learning rate, where the SDE (6) reduces to the gradient flow ODE: $\mathop{}\!\mathrm{d}\bm{\nu}_{t}=-\mathbf{W}^{\top}\mathbf{W}\mathbf{H}\bm{\nu}_{t}\mathop{}\!\mathrm{d}t.$ Noting that $\mathbf{W}^{\top}\mathbf{W}\mathbf{H}=\operatorname{diag}\{\lambda_{1},\lambda_{2},\cdots,\lambda_{M},0,\cdots,0\}$ is diagonal, consequently the ODE is solvable and gives the following expression of the excess risk:
$\mathcal{R}(\bm{\nu}_{t})-\frac{1}{2}\sigma^{2}\eqsim\sum_{j=1}^{M}\lambda_{j}|\theta_{j}^{*}|^{2}e^{-2\lambda_{j}t}+\sum_{j=M+1}^{N}\lambda_{j}|\theta_{j}^{*}|^{2}.$
@@ 4 Functional Scaling Laws
Intuitively, we can view the learning of each eigenfunction as a subtask. Due to the limited model size, student model can at most learn the top-$M$ eigenfunctions.
@@
    Intrinsic-time power law. For each sub-task, the sub-task risk converges exponentially w.r.t. the intrinsic time $t$. However, owing to the power-law structure of $\lambda_{j},\theta_{j}^{*}$, the total multi-task risk exhibits a power-law decay for sufficiently large $M$ due to $\sum_{j=1}^{M}\lambda_{j}(\theta_{j}^{*})^{2}e^{-2\lambda_{j}t}\eqsim\int_{0}^{1}u^{s-1}e^{-2ut}\mathop{}\!\mathrm{d}u\eqsim\frac{1}{t^{s}}$.
@@
    Model-size power law. Approximation error accounts for total risk of the NMN-M unlearned subtasks, which follows a power-law decay due to $\sum_{j=M+1}^{N}\lambda_{j}(\theta_{j}^{*})^{2}\eqsim M^{-s\beta}$ if $N\gg 1$.
5 The Role of Learning Rate Schedules
In this section, we illustrate how our FSL can help understand the influences of LRS on scaling laws. For simplicity, we make the following assumption:
@@ 5 The Role of Learning Rate Schedules
Assume $s\in(0,2-1/\beta]$ and $b(\tau)=B$ for all $\tau\geqslant 0$, and the FSL (7) holds.
When the condition $s\leqslant 2-1/\beta$ holds – indicating a relatively hard task – the FSL simplifies to
$\mathbb{E}[\mathcal{E}(\bm{\nu}_{t})]\eqsim\frac{1}{M^{s\beta}}+\frac{1}{t^{s}}+\frac{\sigma^{2}}{B}\mathcal{A}(\varphi),\quad\text{ with }\quad\mathcal{A}(\varphi)=\int_{0}^{t}\mathcal{K}(t-r)h\varphi(T^{-1}(r))\mathop{}\!\mathrm{d}r,$
where the fit-dependent noise term $e(r)$ is absorbed by the full-batch GD term due to $s\leqslant 2-1/\beta$. Extending to the full range $s\geqslant 0$ is possible but makes the statements and derivations much more involved. We therefore focus on the above case to streamline the exposition.
To connect our continuous‐time analysis with discrete SGD, we study the excess risk after $K=\tau/h$ steps and write $\mathcal{E}_{K}:=\mathbb{E}[\mathcal{E}(\bm{\nu}_{Kh})]$. We then derive scaling laws characterizing $\mathcal{E}_{K}$ as a function of the model size $M$ and total training step $K$, and how the associated scaling exponents vary with the model capacity, characterized by $\beta$. Specifically, two canonical resource-constrained settings are analyzed: (i) the traditional data-limited regime, where the data size $D:=BK$ is fixed; and (ii) the modern compute-limited regime (hoffmann2022training, ), where the compute $C:=MD$ is fixed.

### v3

@@ 4.4 Key Proof Steps and Core Insights
$(2\rho_{-}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\nabla^{2}\mathcal{R}(\mathbf{v})\;\preceq\;\bm{\Sigma}(\mathbf{v})\;\preceq\;(2\rho_{+}\mathcal{E}(\mathbf{v})+\sigma^{2})\,\nabla^{2}\mathcal{R}(\mathbf{v}),$
where $\nabla^{2}\mathcal{R}(\mathbf{v})=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$, and the constants $\rho_{-}$ and $\rho_{+}$ are the same as in Assumption 2.1.
[…]
$\mathbb{E}[|\bm{\xi}(\mathbf{v})^{\top}\bm{n}|^{2}]=\bm{n}^{\top}\bm{\Sigma}(\mathbf{v})\bm{n}\;\eqsim\;\mathcal{R}(\mathbf{v})\,\bm{n}^{\top}\nabla^{2}\mathcal{R}(\mathbf{v})\bm{n},$
where $\bm{n}^{\top}\nabla^{2}\mathcal{R}(\mathbf{v})\bm{n}$ represents the local curvature of the population risk along $\bm{n}$. Hence, the noise energy in each direction is proportional to the product of the population risk and the curvature along that direction. This anisotropic structure of the gradient noise—scaling with the risk and shaped by curvature—has also been reported in prior work [69, 63].
For clarity, in this section, we focus on the case of top-$M$ features, for which the population risk takes the form
$2\mathcal{E}(\mathbf{v})=\sum_{j=1}^{M}\lambda_{j}(v_{j}-\theta^{*}_{j})^{2}+\sum_{j=M+1}^{N}\lambda_{j}|\theta_{j}^{*}|^{2}.$ (13)
For the intrinsic-time SDE (8), each coordinate of $\bm{\nu}(t)$ evolves as
$\mathop{}\!\mathrm{d}\nu_{j}(t)=-\lambda_{j}(\nu_{j}-\theta^{*}_{j})\mathop{}\!\mathrm{d}t+\sqrt{\gamma(t)q_{j}(t)}\mathop{}\!\mathrm{d}B_{j}(t),$
where $q_{j}(t):=\mathrm{\mathbf{e}}_{j}^{\top}\bm{\Sigma}(\bm{\nu}_{t})\mathrm{\mathbf{e}}_{j}$ is the variance of gradient noise along $\mathrm{\mathbf{e}}_{j}$, and $\mathrm{\mathbf{e}}_{j}$ is the $j$-th canonical basis vector for $j\in[M]$. By applying Itô’s formula to $(\nu_{j}-\theta^{*}_{j})^{2}$ and noting $\bm{\nu}(0)=\bm{0}$, we obtain
$\mathbb{E}[(\nu_{j}(t)-\theta_{j}^{*})^{2}]=(0-\theta_{j}^{*})^{2}e^{-2\lambda_{j}t}+\lambda_{j}\int_{0}^{t}e^{-2\lambda_{j}(t-z)}\gamma(z)q_{j}(z)\mathop{}\!\mathrm{d}z.$
Let $\mathcal{E}_{t}=\mathcal{E}(\bm{\nu}_{t})$ and plugging the above equation into (13) gives
$2\mathbb{E}[\mathcal{E}_{t}]$ $=\sum_{j=1}^{M}\lambda_{j}|\theta_{j}^{*}|^{2}e^{-2\lambda_{j}t}+\sum_{j=1}^{M}\lambda_{j}^{2}\int_{0}^{t}e^{-2\lambda_{j}(t-z)}\gamma(z)\mathbb{E}[q_{j}(z)]\mathop{}\!\mathrm{d}z+\sum_{j=M+1}^{N}\lambda_{j}|\theta_{j}^{*}|^{2}.$ (14)
By Lemma 4.8 and noting $\nabla^{2}\mathcal{R}(\mathbf{v})=\operatorname{diag}(\lambda_{1},\dots,\lambda_{M})$, we have
$q_{j}=\mathrm{\mathbf{e}}_{j}^{\top}\bm{\Sigma}(\bm{\nu}_{t})\mathrm{\mathbf{e}}_{j}\eqsim\lambda_{j}\mathcal{R}(\bm{\nu}_{t})=\lambda_{j}(\mathcal{E}(\bm{\nu}_{t})+\sigma^{2}/2).$
Let $\delta_{M}=\sum_{j=M+1}^{N}\lambda_{j}|\theta_{j}^{*}|^{2}$, $e_{M}(t)=\sum_{j=1}^{M}\lambda_{j}|\theta_{j}^{*}|^{2}e^{-2\lambda_{j}t}$, and $\mathcal{K}_{M}(t)=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}$. Plugging them back into (14) gives the following Volterra equation:
$\mathbb{E}[\mathcal{E}_{t}]\eqsim\delta_{M}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-z)\gamma(z)(\mathbb{E}[\mathcal{E}_{t}]+\sigma^{2})\mathop{}\!\mathrm{d}z.$ (15)
The above equation characterizes the expected loss dynamics of SGD under a general spectrum and has been derived in prior works such as [48, 49, 30, 50, 51]. Our key observation is that, under the power-law assumptions on $\{\theta_{j}^{*}\}_{j}$ and $\{\lambda_{j}\}_{j}$ (Assumptions 2.3 and 2.5), the solution to (15) admits a sharp asymptotic characterization, providing explicit upper and lower bounds that precisely capture its scaling behavior.
Let $f(t):=\mathbb{E}[\mathcal{E}_{t}]$, $g(t):=\delta_{M}+e_{M}(t)+\sigma^{2}\int_{0}^{t}\mathcal{K}_{M}(t-z)\gamma(z)\mathop{}\!\mathrm{d}z$, and define the linear operator
$\mathcal{T}f(t)=\int_{0}^{t}\mathcal{K}_{M}(t-z)\gamma(z)f(z)\mathop{}\!\mathrm{d}z.$
Then, the Volterra equation (15) can be expressed in the compact form $f=g+\mathcal{T}f.$ Formally, its solution can be expanded as an infinite series:
$f=(\mathcal{I}-\mathcal{T})^{-1}g=g+\mathcal{T}g+\mathcal{T}^{2}g+\mathcal{T}^{3}g+\cdots.$ (16)
The key observation is that, under Assumptions 2.3 and 2.5, the higher-order terms $\mathcal{T}^{k}g$ for $k\!\geqslant\!2$ can be well controlled by the first-order term $\mathcal{T}g$. This is due to the scale invariance of the forgetting kernel:
Lemma 4.9 (Scale invariance).
It holds for any $t\lesssim M^{\beta}$ that $\tfrac{\mathcal{K}_{M}(t/2)}{\mathcal{K}_{M}(t)}\simeq 1$.
Proof.
The result follows from the fact that $\mathcal{K}_{M}$ exhibits a power-law decay for $t\!\lesssim\!M^{\beta}$ (see (20)). Consequently, $\mathcal{K}_{M}(t/2)\eqsim(t/2)^{-(2-1/\beta)}\eqsim t^{-(2-1/\beta)}\eqsim\mathcal{K}_{M}(t),$ which establishes the claim. ∎
Corollary 4.10 (Subconvolution property).
For any $t\lesssim M^{\beta}$, it holds $\mathcal{K}_{M}*\mathcal{K}_{M}(t)\lesssim\mathcal{K}_{M}(t)$.
Proof.
Noting $\mathcal{K}_{M}$ is non-increasing and integrable and applying Lemma 4.9, we obtain
$(\mathcal{K}_{M}*\mathcal{K}_{M})(t)$ $=\int_{0}^{t/2}\mathcal{K}_{M}(t-z)\mathcal{K}_{M}(z)\mathop{}\!\mathrm{d}z+\int_{t/2}^{t}\mathcal{K}_{M}(t-z)\mathcal{K}_{M}(z)\mathop{}\!\mathrm{d}z$
$\leqslant 2\,\mathcal{K}_{M}(t/2)\!\int_{0}^{t/2}\!\mathcal{K}_{M}(z)\mathop{}\!\mathrm{d}z\;\lesssim\;\mathcal{K}_{M}(t/2)\;\lesssim\;\mathcal{K}_{M}(t),$
which proves the claim. ∎
Let $\|\gamma\|_{\infty}=\sup_{t\geqslant 0}\gamma(t)$. By Corollary 4.10, it holds for any $t\lesssim M^{\beta}$ that
$\mathcal{T}^{2}g(t)$ $\leqslant\|\gamma\|_{\infty}\int_{0}^{t}\mathcal{K}_{M}*\mathcal{K}_{M}(t-z)\gamma(z)g(z)\mathop{}\!\mathrm{d}z$
$\lesssim\|\gamma\|_{\infty}\int_{0}^{t}\mathcal{K}_{M}(t-z)\gamma(z)g(z)\mathop{}\!\mathrm{d}z=\|\gamma\|_{\infty}\mathcal{T}g(t).$
When $\|\gamma\|_{\infty}$ is sufficiently small, there exists a constant $0<c<1$ such that $\mathcal{T}^{k}g(t)\leqslant c^{k-1}\mathcal{T}g(t)$ holds for any $0\leqslant t\lesssim M^{\beta}$. Hence,
$f(t)\leqslant g(t)+\mathcal{T}g(t)+\sum_{k=2}^{\infty}c^{k-1}\mathcal{T}g(t)\lesssim g(t)+\mathcal{T}g(t)+\tfrac{c}{1-c}\mathcal{T}g(t).$
Combining $f(t)\geqslant g(t)+\mathcal{T}g(t)$ with the above upper bound, we conclude $f(t)\eqsim g(t)+\mathcal{T}g(t)$, i.e., we have
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\tfrac{1}{2}\sigma^{2}\;\eqsim\;\delta_{M}+e_{M}(t)+\int_{0}^{t}\mathcal{K}_{M}(t-z)\,[e_{M}(z)+\sigma^{2}]\,\gamma(z)\,\mathop{}\!\mathrm{d}z.$ (17)
The Emergence of Power-Law Scaling.
@@
We now illustrate how power-law scaling arises in our setting from a multi-task learning viewpoint. Intuitively, we can view the learning of each eigenfunction as a sub-task. Due to the limited model size, student model can at most learn the top-$M$ eigenfunctions. Within this view, our FSL framework reveals three distinct manifestations of power-law behavior:
@@
    Approximation error. Approximation error accounts for total risk of the NMN-M unlearned sub-tasks, which follows a power-law decay due to
    $\delta_{M}=\sum_{j=M+1}^{N}\lambda_{j}|\theta_{j}^{*}|^{2}\eqsim M^{-s\beta},\text{ if }N-M\gtrsim M.$ (18)
@@
    Signal learning. For each sub-task, the sub-task risk converges exponentially w.r.t. the intrinsic time $t$. However, owing to the power-law structure of $\{\lambda_{j}\},\{\theta_{j}^{*}\}$, the total multi-task risk exhibits a power-law decay for sufficiently large $M$ due to
    $e_{M}(t)=\sum_{j=1}^{M}\lambda_{j}|\theta_{j}^{*}|^{2}e^{-2\lambda_{j}t}\eqsim\int_{M^{-\beta}}^{1}u^{s-1}e^{-2ut}\mathop{}\!\mathrm{d}u\eqsim\frac{1}{t^{s}},\text{ if }1\lesssim t\lesssim M^{\beta}.$ (19)
  • (iii)
    Noise forgetting. Analogously, for the forgetting kernel, we also have
    $\mathcal{K}_{M}(t)=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}\;\eqsim\;\int_{M^{-\beta}}^{1}z^{\,1-\frac{1}{\beta}}e^{-2zt}\mathop{}\!\mathrm{d}z\;\eqsim\;t^{-(2-1/\beta)},\text{ if }1\lesssim t\lesssim M^{\beta}.$ (20)
Remark 4.11 (Task Accumulation Effect).
The above derivation reveals that, although a single task may not exhibit an exact power-law scaling with respect to intrinsic time, the collective accumulation of many tasks gives rise to such scaling behavior. As the number of tasks increases, the power-law regime progressively extends, and in the idealized limit $M\to\infty$, it spans the entire training process. In our setting, each sub-task naturally follows an exponential learning curve $\exp(-\lambda t)$, yet the underlying mechanism appears more universal. We anticipate that similar power-law behavior may arise even when individual task dynamics deviate from this form, which we leave for future investigation.
5 Learning Rate Schedules Impact Scaling Efficiency
Having established the general FSL, we now instantiate it under three representative LRSs—constant, exponential decay, and warmup–stable–decay (WSD)—to examine how schedule design influences scaling efficiency. All proofs can be found in Appendix D. For clarity, we make:
@@ 5 Learning Rate Schedules Impact Scaling Efficiency
Assume constant label noise $\sigma^{2}\gtrsim 1$ and batch size $b(\tau)=B$ for all $\tau\geqslant 0$.
Under this assumption, given a physical-time LRS function $\varphi(\cdot)$, Theorem 4.4 implies that the FSL for $t\gtrsim 1$ simplifies to
$\mathbb{E}[\mathcal{R}(\bm{\nu}_{t})]-\tfrac{1}{2}\sigma^{2}$ $\eqsim M^{-s\beta}+e_{M}(t)+\frac{\sigma^{2}}{B}\int_{0}^{t}\mathcal{K}_{M}(t-r)\varphi(T^{-1}(r))\mathop{}\!\mathrm{d}r.$
Let $\mathcal{E}_{K}=\mathbb{E}[\mathcal{R}(\bm{\nu}_{Kh})]-\tfrac{1}{2}\sigma^{2}$ denote the expected excess risk after $K$ training steps. For each LRS, we derive concrete scaling laws describing how $\mathcal{E}_{K}$ scales with the model size $M$, the total step count $K$, as well as the LRS’s hyperparameters. We then reinterpret these results from a resource-allocation perspective by optimizing under two canonical constraints: (i) the data-limited regime, where the total data size $D:=BK$ is fixed; and (ii) the compute-limited regime [21], where the total compute $C:=MD$ is fixed. For each regime, we further examine how the optimally tuned hyperparameters (e.g., the peak learning rate) should scale with increasing available resources.
Finally, for clarity, we distinguish between two task regimes: an easy-learning regime, where $s\geqslant 1-1/\beta$, and a hard-learning regime, where $s<1-1/\beta$.